In a dog show there are 3 times as many collies as boxers. There are also 3 poodles for every 5 collies. What is the smallest possible number of dogs that could be in the dog show?
Thank You!

Answer by stanbon(75887)

In a dog show there are 3 times as many collies as boxers. There are also 3 poodles for every 5 collies. What is the smallest possible number of dogs that could be in the dog show?
-----------
# of boxers = "x"
# of collies = "3x"
# of poodles = (3/5)(3x) = (9/5)x
-------------
Let x = 5 (# of boxers)
Then 3x = 15 (# of collies)
And (9/5)x = 9 (# of poodles)
-------------
Cheers,
Stan H.